OpenStudy (anonymous):

The graph below shows the distance (y) in kilometers of two cars from their destination at different times (x) in minutes. The table shows the values plotted on the graph: Time (in minutes) x 0 10 20 30 40 50 60 ------------------------------------------ Car 1 (in km) y 60 50 40 30 20 10 0 ------------------------------------------ Car 2 (in km) y 60 49.02438 40.05648 32.72904 26.742 21.8502 17.85318 ------------------------------------------

OpenStudy (anonymous):

Part D: What is the average rate of change from x = 20 to x = 30 for the function representing the motion for car 2? What does the value of this average rate of change represent? (4 points)

OpenStudy (anonymous):

I need help writing the function for Car 2, the rest I can do myself. PLEASE HELP! WILL GIVE A MEDAL!

OpenStudy (anonymous):

@iGreen @SolomonZelman

OpenStudy (igreen):

Sorry this is confusing..could you take a picture of it?

OpenStudy (anonymous):

Ya there's a graph for it

OpenStudy (anonymous):

Sorry I didn't post that earlier :|

OpenStudy (igreen):

Hmm..the y-intercept is 60.

OpenStudy (anonymous):

yes this is what I got so far, because I think it should be an exponential function: f(x) = 60(b)^x

OpenStudy (igreen):

Yeah it's definitely exponential..

OpenStudy (anonymous):

I just don't what b is. I know b is the base of the exponential function and it models the change that is happening.

OpenStudy (igreen):

Yes..

OpenStudy (igreen):

Okay it's gonna be.. f(x) = 60(0.98^x)

OpenStudy (anonymous):

Can you explain how you got 0.98?

OpenStudy (igreen):

Slope from (0, 60) and (2, 40) is approximately 0.98. (Using the slope formula)

OpenStudy (igreen):

Graph it here: https://www.desmos.com/calculator

OpenStudy (igreen):

You will see the exact same graph.

OpenStudy (igreen):

Need help with Part D?

OpenStudy (anonymous):

No, but thx with everything!!!!!!!!!!

OpenStudy (anonymous):

Will u be here if I need help with another question?

OpenStudy (igreen):

Yep.

OpenStudy (anonymous):

THX.... I luv u! I actually don't really luv u(like that) but u know.........

OpenStudy (anonymous):

as a friend

OpenStudy (anonymous):

-.-

OpenStudy (igreen):

Lel

OpenStudy (igreen):

Btw, do you do FLVS?

OpenStudy (igreen):

I'm pretty sure you do..

OpenStudy (anonymous):

Yup

OpenStudy (igreen):

'Cause I remember this stuff..

OpenStudy (igreen):

Algebra I?

OpenStudy (anonymous):

Did u take FLVS or are u taking it?

OpenStudy (igreen):

I'm still taking FLVS, that's my school. But I've already finished Algebra I..

OpenStudy (anonymous):

are u still there @iGreen , because I think I need help with the average rate of change, since, I got a negative decimal

OpenStudy (igreen):

Okay.

OpenStudy (anonymous):

\[\frac{ f(b)-f(a) }{ b-a }= \frac{ 40.05-32.72 }{ 20-30 }= \frac{ 7.33 }{ -10 }\]

OpenStudy (anonymous):

that doesn't even make sense

OpenStudy (anonymous):

so maybe I did something wrong?..............

OpenStudy (igreen):

What is the average rate of change from x = 20 to x = 30 for the function representing the motion for car 2? This is the formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) Points: \((20, 40), (30, 32.73)\) \(y_2 = 32.73\) \(y_1 = 40\) \(x_2 = 30\) \(x_1 = 20\) Plug it in the formula: \(m = \dfrac{32.73-40}{30-20}\) Simplify: \(m = \dfrac{-7.27}{10}\) \(m = -0.727\)

OpenStudy (igreen):

It's supposed to be negative..because the slope is also negative..

OpenStudy (igreen):

@Chocolate1031

OpenStudy (anonymous):

oh I see, I forgot that formula.......... I LIKE ALWAYS DO, MY DAD KEEPS ON TELLING ME

OpenStudy (anonymous):

I am wierd

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Lol